if a+b+c=0 and a^2+b^2+c^2=k(a^2-bc) then value of k is

Grade:12

1 Answers

divyam gupta
5 years ago
(a+b+c)^2=a^2+b^2+c^2+2ab+2bc+2ca=0
a^2+b^2+c^2=2(-a(b+c)-bc)
from given condition:
a+b+c=0
therfore:b+c=-a
use above condition to get
a^2+b^2+c^2=2(a^2-bc)
therfore,k=2